# NVRC: Neural Video Representation Compression

### v1

@@ 3.2 Network layer parameter coding
Unlike the feature grids in INR models, the network parameters, such as the weights in linear and convolutional layers, are difficult to compress as there is no spatial or temporal correlation between parameters. Existing works typically use simple entropy models for encoding these parameters [11, 25, 14, 38, 28, 58].
@@
2D block partitioning. In the proposed NVRC framework, similar to feature grids, network layer parameters are also partitioned into groups prior to quantization and coding. Here we assume that the encoding of the parameters belonging to the same input/output features/channels could benefit from sharing quantization and entropy coding parameters. For example, if an input feature/channel is zero, then the corresponding group of parameters are likely to be zeros as well. In NVRC, the quantization and entropy coding models for network parameters are designed based on this assumption, and aim to share the quantization and entropy models between parameters in the same row/column. Since there are parameters with different numbers of dimensions, e.g., 2D for linear layer weights and 4/5D for convolution weights, all layer parameters in NVRC are first reshaped into 2D tensors, where the parameters in a row correspond to the weight from the same output feature/channel, and the parameters in a column are the weights for the same input feature/channel. While existing works [11, 25] can be directly employed on partitioned parameter tensors by rows or columns, and applied per-row or per-column entropy parameters for coding, this may not be the best solution, because (i) the partitioning axis needs to be decided, (ii) the coding could benefit from sharing quantization and entropy parameters across both rows and columns. Therefore, in NVRC, the tensors are partitioned according to both axes at the same time, and the quantization and entropy parameters are learned in both axes and mixed during coding. We noticed that existing work has utilized quantization parameters on both input and output channels in different contexts [13]. Here, entropy parameters are also utilized, and both the quantization and entropy parameters are further compressed.
@@
Quantization. For the weights of a layer, $\omega^{layer}$, from $\theta^{layer}$, $\omega^{layer}\in\mathbb{R}^{C_{out}\times C_{in}}$, the quantization scales $\delta^{layer}\in\mathbb{R}^{C_{out}\times C_{in}}$, is combined by two vectors of scales $\delta^{layer}_{out}\in\mathbb{R}^{C_{out}}$ and $\delta^{layer}_{in}\in\mathbb{R}^{C_{in}}$, such that:
$\delta^{layer}[i,j]=\delta^{layer}_{out}[i]\times\delta^{layer}_{in}[j],$ (5)
@@
where $0\leq i<C_{out}$ and $0\leq j<C_{in}$. In practice, only the logarithms of $\delta^{layer}_{out}$ and $\delta^{layer}_{in}$ are stored, and quantization is performed similarly to the grid parameters (Section 3.1).
@@
Dual-axis conditional Gaussian model. In NVRC, a dual-axis conditional Gaussian model is used for coding the network layer parameters. Similar to the quantization parameters mentioned above, the means $\mu^{layer}$ and the scales $\sigma^{layer}$, are represented in two per-axis parameter vectors, i.e. $\mu^{layer}_{out}$, $\mu^{layer}_{in}$, $\sigma^{layer}_{out}$ and $\sigma^{layer}_{in}$, and they are both from $\hat{\phi}_{em}$.
The combined means $\mu^{layer}$ and scales $\sigma^{layer}$ are obtained by
$\mu^{layer}[i,j]=\mu^{layer}_{out}[i]\times\sigma^{layer}_{in}[j]+\mu^{layer}_{in}[j],$ (6)
@@
$\sigma^{layer}[i,j]=\sigma^{layer}_{out}[i]\times\sigma^{layer}_{in}[j],$ (7)
where $0\leq i<C_{out},0\leq j<C_{in}$. Like the quantization parameters $\delta^{layer}$, only the per-axis means $\mu^{layer}_{out}$, $\mu^{layer}_{in}$, and the logarithms of the per-axis scales $\sigma^{layer}_{out}$, $\sigma^{layer}_{in}$ are stored. Finally, the means and scales here will also (as for feature grids) be scaled by $\delta^{layer}$, before being utilized for coding $\hat{\omega}^{layer}$.
@@ 3.4 Rate-distortion optimization
Alternating optimization. In existing INR-based video representations and compression methods [11, 25], the distortion loss is minimized iteratively with sampling batches of frames, patches or pixels. To introduce the entropy regularization, this process has been extended [24, 58], where the rate loss is also calculated in each training step, similar to other learning-based video compression methods [35, 29]. However, in the INR-based video compression, the training is the process for over-fitting the network to a video sequence, in which the samples of each steps are from the same sequence, and the code, i.e., the INR model parameters, is the same set of parameters for all steps. Thus, it is not necessary to update the rate term in every step, especially when a significant amount of computation or memory is needed for this due to the use of entropy models. In NVRC, a more efficient training process is used, where the rate $R$ and distortion $D$ are optimized alternately. In every $K+1$ steps, the $D$ is minimized in the first $K$ steps, and where $R$ is minimized at the $K+1$-th step. Empirically, the rate loss is also scaled by $K$ to keep the rate roughly the same. Note that, the quantization step is still applied on each step, and skipping the entropy model is only possible when the quantization parameters are separated from the entropy model.
Two-stage training. Similar to some existing works [11, 25, 24], NVRC is also trained in two stages.
In Stage 1, to optimize $L$, the non-differentiable quantization operation needs to be emulated through a differentiable approximation during training. Recent work [24] has shown that a soft-rounding operation with an additive Kumaraswamy noise can be used to replace quantization for neural representation training. While in [24], this is applied only to feature grids, we extend this idea and apply it to both feature grids and network parameters in the first stage of training. Compared to [24], soft-rounding with higher temperature (0.5 to 0.3) is used in NVRC, as the original, low temperature (e.g. 0.3 to 0.1 in [24]) for both feature grids and network parameters will lead to training difficulty due to the large variance of the gradients.
In the second stage, instead of using soft-rounding, following [25, 26], Quant-Noise [48] is used to fine-tune the neural representation, as we empirically found that Quant-Noise remains stable with different hyper-parameter settings and is suitable for high quantization levels.
@@ 4.1 Experiment Configuration
Evaluation database. To evaluate the performance of the proposed NVRC framework, we conducted experiments on the UVG [40] dataset, which contain seven video sequences content with a $1920\times 1080$ resolution and 300/600 frames. Results with additional datasets will be provided in the later version of the paper.
Implementation details. NVRC is a new framework focusing on INR model compression, which can be integrated with any typical INR-based models. To test its effectiveness, we employed one of the latest INR network architectures, HiNeRV [25], and integrated it into our NVRC framework. This INR model has been reported to provide competitive performance compared to many standard and end-to-end codecs for the video compression task. Minor adjustments have been made on top of HiNeRV in terms of the network structure and the training configuration (see Supplementary for details) - rate points are now obtained by both turning the scale of the neural representation and the $\lambda$ value. The model is trained with 360 epochs in the first stage, and 30 epochs in the second stage.
Benchmark methods. Conventional codecs, x265 [1] with the veryslow configuration, HM-18.0 [45] and VTM-20.0 [9] with the Random Access configuration, are used for benchmarking, together with two recent learned video codecs, DCVC-HEM [30], DCVC-DC [31]. Three state-of-the-art INR-based codecs, including the original HiNeRV [25], C3 [24] and HNeRV-Boost [58] have also been included in this experiment. All results are produced by the open source implementations, except for those of C3, which are obtained from the original paper.
Evaluation methods. The evaluation was performed in the RGB color space (for comparing both conventional codecs and learning-based methods) with the BT.601 color conversion, and in the original YUV420 color space (for comparing both conventional methods and the learning-based methods that support this feature in their public implementations). PSNR and MS-SSIM are used here to assess video quality, based on which Bjøntegaard Delta Rate figures are calculated against each benchmark codec.
@@ 4.2 Results and discussion
Table 1: BD-rate results on the UVG dataset.
@@
RGB 4:4:4 PSNR -74.02% -51.00% -24.34% -41.30% -32.05% -50.73% -67.93% -66.78%
MS-SSIM -80.79% -67.61% -50.08% -7.91% -12.58% -44.69% - -78.21%
YUV 4:2:0 PSNR -62.71% -34.83% -1.03% - -62.28% - - -
MS-SSIM -59.49% -38.45% -15.38% - -70.23% - - -
Figure 3: Average rate quality curves of various tested codecs on the UVG datasets.
Figure 3 and Table 1 provide the results for the proposed NVRC model and the benchmark methods. It can be observed that when tested in the RGB 4:4:4 color space (as in many learning-based works), NVRC significantly outperforms the original HiNeRV model, with an average coding gain of 50.73%, measured by PSNR. Similar improvement has also been achieved against other INR-based methods including HNeRV-Boost and C3. Moreover, NVRC also offers better performance compared to latest MPEG standard codec VVC VTM (Random-Access), with a 24% average coding gain based on PSNR. To the best of our knowledge, it is the first INR-based video codec outperforming VTM. Compared to state-of-the-art learned video coding methods, NVRC also exhibits superior performance, beating DCVC-DC and DCVC-HEM. When evaluated in the YUV 4:2:0 color space, NVRC still offers superior performance as for RGB 4:4:4 color space, outperforming most benchmarked methods based on PSNR and MS-SSIM. It should be also noted that INR-based video codecs do not require offline training on large-scale datasets, whereas other learning-based methods do. Qualitative results are provided in Figure 1 in terms of visual comparison between the content reconstructed by NVRC and HiNeRV.
@@ 4.3 Computational complexity
Table 2: Complexity results of NVRC. Encoding and decoding FPS are measured by the number of training steps/evaluation steps per second performed by the INR model with 1080p inputs/outputs. The model compression MACs and encoding/decoding time are measured by the steps for performing quantization and entropy coding. The complexity figures are calculated based on NVIDIA 4090 GPU with FP16.
@@
The complexity figures of NVRC are provide in Table 2. When compared to the original HiNeRV, the proposed method (with HiNeRV as its INR network) is associated with increased computational complexity. However, the MACs is still significantly lower than that of other learning-based video codecs (e.g., DCVC-DC [31]), which allows faster decoding. It should be noted that the complexity figures shown here are obtained based on research source code that has not been optimized for latency. The actual latency of INR and entropy coding can be further reduced by (1) optimizing the implementation of the INR and entropy models, (2) performing lower precision computation, and (3) implementing parallel decoding between different resolution feature grids.
@@ 4.4 Ablation study
Table 3: Ablation studies on the UVG dataset [40].
Model NVRC (V1) (V2) (V3) (V4) (V5)
BD-Rate (in %, PSNR) 0.00 12.85 9.79 21.84 22.12 1.05
BD-Rate (in %, MS-SSIM) 0.00 12.40 9.09 21.32 21.77 1.40
To evaluate the contribution of the main components in NVRC, an ablation study was performed based on the UVG dataset, using the configurations in Section 4.1, but with two sequences (Jockey and ReadySetGo from the UVG dataset [40]) and four rate points for each variant. Results with additional configurations will be provided in the later version of the paper.
Alternative entropy model settings. We compared different combinations of entropy models for encoding feature grids $\theta_{grid}$ and network parameters $\theta_{layer}$: Context model + dual-axis conditional Gaussian model (Default setting in NRVC), (V1) Context model + per-tensor conditional Gaussian model, (V2.) per-tensor conditional Gaussian model + dual-axis conditional Gaussian model, (V3) per-tensor conditional Gaussian model + per-tensor conditional Gaussian model.
@@
Learned quantization steps. We also compared the use of learned quantization steps $\phi_{quant}$ and $\psi_{quant}$ (Default setting in NRVC) and (V5), a new variant with fixed quantization steps for grids, where the log-step size is set to $-4$.
@@
We also performed simple evaluation on other alternatives: (1) Removing the alternating optimization, which has almost no impact on the rate-distortion performance, but at a cost of significant computational overhead; (2) Removing the learnable quantization step size for the network layer parameters, which leads to a significant drop in the rate-distortion performance.

### v2

@@ 3.2 Network layer parameter coding
Unlike the feature grids in INR models, the network parameters, such as the weights in linear and convolutional layers, are difficult to compress as there is no spatial or temporal correlation between parameters. Existing works typically use simple entropy models for encoding these parameters chen2021nerv; kwan2023hinerv; gomes2023video; maiya2023nirvana; leguay2024cool; zhang2024boosting.
@@
2D block partitioning. In the proposed NVRC framework, similar to feature grids, network layer parameters are also partitioned into groups prior to quantization and coding. Here we assume that the encoding of the parameters from the same input/output features/channels could benefit from sharing quantization and entropy coding parameters. For example, if an input feature/channel is zero, then the corresponding group of parameters are likely to be zeros as well. In NVRC, the quantization and entropy coding models for network parameters are designed based on this assumption, and aim to share the quantization and entropy models between parameters in the same row/column. Since there are parameters with different numbers of dimensions, e.g., 2D for linear layer weights and 4/5D for convolution weights, all layer parameters in NVRC are first reshaped into 2D tensors, where the parameters in a row correspond to the weight from the same output feature/channel, and the parameters in a column are the weights for the same input feature/channel. While existing works chen2021nerv; kwan2023hinerv can be directly employed on partitioned parameter tensors by rows or columns, and applied per-row or per-column entropy parameters for coding, this may not be the best solution, because (i) the partitioning axis needs to be decided, (ii) the coding could benefit from sharing quantization and entropy parameters across both rows and columns. Therefore, in NVRC, the tensors are partitioned according to both axes at the same time, and the quantization and entropy parameters are learned in both axes and mixed during coding. We noticed that existing work has utilized quantization parameters on both input and output channels in different contexts finkelstein22qft. Here, entropy parameters are also utilized, and both the quantization and entropy parameters are further compressed.
@@
Quantization. For the weights of a layer, $\omega_{layer}$, from $\theta_{layer}$, $\omega_{layer}\in\mathbb{R}^{C_{out}\times C_{in}}$, the quantization scales $\delta_{layer}\in\mathbb{R}^{C_{out}\times C_{in}}$, is combined by two vectors of scales $\delta_{layer,out}\in\mathbb{R}^{C_{out}}$ and $\delta_{layer,in}\in\mathbb{R}^{C_{in}}$, such that:
$\delta_{layer}[i,j]=\delta_{layer,out}[i]\times\delta_{layer,in}[j],$ (5)
@@
where $0\leq i<C_{out}$ and $0\leq j<C_{in}$. In practice, only the logarithms of $\delta_{layer,out}$ and $\delta_{layer,in}$ are stored, and quantization is performed similarly to the grid parameters (Section 3.1).
@@
Dual-axis conditional Gaussian model. In NVRC, a dual-axis conditional Gaussian model is used for coding the network layer parameters. Similar to the quantization parameters mentioned above, the means $\mu_{layer}$ and the scales $\sigma_{layer}$, are represented in two per-axis parameter vectors, i.e. $\mu_{layer,out}$, $\mu_{layer,in}$, $\sigma_{layer,out}$ and $\sigma_{layer,in}$, and they are both from $\hat{\phi}_{em}$.
The combined means $\mu_{layer}$ and scales $\sigma_{layer}$ are obtained by
$\mu_{layer}[i,j]=\mu_{layer,out}[i]\times\sigma_{layer,in}[j]+\mu_{layer,in}[j],$ (6)
@@
$\sigma_{layer}[i,j]=\sigma_{layer,out}[i]\times\sigma_{layer,in}[j],$ (7)
where $0\leq i<C_{out},0\leq j<C_{in}$. Like the quantization parameters $\delta_{layer}$, only the per-axis means $\mu_{layer,out}$, $\mu_{layer,in}$, and the logarithms of the per-axis scales $\sigma_{layer,out}$, $\sigma_{layer,in}$ are stored. Finally, the means and scales here will also (as for feature grids) be scaled by $\delta_{layer}$, before being utilized for coding $\hat{\omega}_{layer}$.
@@ 3.4 Rate-distortion optimization
Alternating optimization. In existing INR-based video representations and compression methods chen2021nerv; kwan2023hinerv, the distortion loss is minimized iteratively with sampling batches of frames, patches or pixels. To introduce the entropy regularization, this process has been extended kim2023c3; zhang2024boosting, where the rate loss is also calculated in each training step, similar to other learning-based video compression methods lu2019dvc; li2021deep. However, in the INR-based video compression, the training is the process for over-fitting the network to a video sequence, in which the samples of each steps are from the same sequence, and the code, i.e., the INR model parameters, is the same set of parameters for all steps. Thus, it is not necessary to update the rate term in every step, especially when a significant amount of computation or memory is needed for this due to the use of entropy models. In NVRC, a more efficient training process is used, where the rate $R$ and distortion $D$ are optimized alternately. In every $K+1$ steps, the $D$ is minimized in the first $K$ steps, and where $R$ is minimized at the $K+1$-th step. Empirically, the rate loss is also scaled by $K$ to keep the rate roughly the same. Note that, the quantization step is still applied on each step, and skipping the entropy model is only possible when the quantization parameters are separated from the entropy model.
Two-stage training. Similar to some existing works chen2021nerv; kwan2023hinerv; kim2023c3, NVRC is also trained in two stages.
In Stage 1, to optimize $L$, the non-differentiable quantization operation needs to be emulated through a differentiable approximation during training. Recent work kim2023c3 has shown that a soft-rounding operation with an additive Kumaraswamy noise can be used to replace quantization for neural representation training. While in kim2023c3, this is applied only to feature grids, we extend this idea and apply it to both feature grids and network parameters in the first stage of training. Compared to kim2023c3, soft-rounding with higher temperature (0.5 to 0.3) is used in NVRC, as the original, low temperature (e.g. 0.3 to 0.1 in kim2023c3) for both feature grids and network parameters will lead to training difficulty due to the large variance of the gradients.
In the second stage, instead of using soft-rounding, following kwan2023hinerv; kwan2024immersive, Quant-Noise fan2020training is used to fine-tune the neural representation, as we empirically found that Quant-Noise remains stable with different hyper-parameter settings and is suitable for high quantization levels.
@@ 4.1 Experiment Configuration
Evaluation database. To evaluate the performance of the proposed NVRC framework, we conducted experiments on the UVG mercat2020uvg and MCL-JCV wang2016mcl dataset. The UVG dataset includes 7 video sequences with 300/600 frames, while the MCL-JCV dataset consists of 30 video clips with 120-150 frames. All sequences are compressed at their original resolution in this experiment. We also provide the result of JVET-CTC dataset Class B boyce2018jvet in the Appendix.
Implementation details. NVRC is a new framework focusing on INR model compression, which can be integrated with any typical INR-based models. To test its effectiveness, we employed one of the latest INR network architectures, HiNeRV kwan2023hinerv, and integrated it into our NVRC framework. This INR model has been reported to provide competitive performance compared to many standard and end-to-end codecs for the video compression task. Minor adjustments have been made on top of HiNeRV in terms of the network structure and the training configuration (see Appendix for details) - rate points are now obtained by both turning the scale of the neural representation and the $\lambda$ value. The model is trained for 360 or 720 epochs in the first stage and 30 or 60 epochs in the second stage, depending on the UVG mercat2020uvg and MCL-JCV wang2016mcl datasets, due to the differing lengths of the sequences.
Benchmark methods. Conventional codecs, x265 x265 with the veryslow configuration, HM-18.0 hm and VTM-20.0 vtm with the Random Access configuration, are used for benchmarking, together with two recent learned video codecs, DCVC-HEM li2022hybrid, DCVC-DC li2023neural. Three state-of-the-art INR-based codecs, including the original HiNeRV kwan2023hinerv, C3 kim2023c3 and HNeRV-Boost zhang2024boosting have also been included in this experiment. All results are produced by the open source implementations.
Evaluation methods. The evaluation was performed in the RGB color space (for comparing both conventional codecs and learning-based methods) with the BT.601 color conversion, and in the original YUV420 color space (for comparing both conventional methods and the learning-based methods that support this feature in their public implementations). PSNR (RGB/YUV 6:1:1) and MS-SSIM (RGB/Y) are used here to assess video quality, based on which Bjøntegaard Delta Rate figures are calculated against each benchmark codec.
@@ 4.2 Results and discussion
Table 1: BD-rate results on the UVG dataset mercat2020uvg.
@@
RGB 4:4:4 PSNR -73.74% -50.38% -23.42% -40.57% -31.20% -50.16% -66.86% -66.45%
MS-SSIM -80.65% -67.38% -49.75% -6.97% -11.75% -44.27% -76.59% -78.01%
YUV 4:2:0 PSNR -66.89% -42.50% -12.96% - -33.98% - - -
MS-SSIM -59.38% -38.20% -15.04% - -39.12% - - -
Table 2: BD-rate results on the MCL-JCV dataset wang2016mcl.
RGB 4:4:4 PSNR -51.61% -13.88% 36.91% -2.97% 13.40% -31.69% -42.23% -59.80%
MS-SSIM -66.83% -41.01% -6.39% -21.64% 33.07% -41.62% -49.23% -83.36%
YUV 4:2:0 PSNR -49.02% -13.19% 40.80% - 3.75% - - -
MS-SSIM -43.00% -12.61% 33.09% - -9.89% - - -
Figure 3: Average rate quality curves of various tested codecs on the UVG dataset mercat2020uvg.
Figure 4: Average rate quality curves of various tested codecs on the MCL-JCV dataset wang2016mcl.
Figure 3-4 and Table 1-2 provide the results for the proposed NVRC model and the benchmark methods. It can be observed that when tested in the RGB 4:4:4 color space (as in many learning-based works), NVRC significantly outperforms the original HiNeRV model kwan2023hinerv, with an average coding gain of 50.16%, measured by PSNR. Similar improvement has also been achieved against other INR-based methods including HNeRV-Boost zhang2024boosting and C3 kim2023c3. Moreover, NVRC also offers better performance compared to latest MPEG standard codec VVC VTM (Random-Access) vtm on the UVG dataset mercat2020uvg, with a 23.4% average coding gain based on PSNR. To the best of our knowledge, it is the first INR-based video codec outperforming VTM. Compared to state-of-the-art learned video coding methods, NVRC also exhibits superior or comparable performance to DCVC-HEM li2022hybrid and DCVC-DC li2023neural on the UVG mercat2020uvg and MCL-JCV wang2016mcl datasets, respectively.. When evaluated in the YUV 4:2:0 color space, NVRC still offers superior performance as for RGB 4:4:4 color space, outperforming most benchmarked methods based on PSNR and MS-SSIM. It should be also noted that INR-based video codecs do not require offline training on large-scale datasets, whereas other learning-based methods do. Qualitative results are provided in Figure 1 in terms of visual comparison between the content reconstructed by NVRC and HiNeRV.
@@ 4.3 Computational complexity
Table 3: Complexity results of NVRC with the UVG dataset mercat2020uvg. Encoding and decoding FPS are measured by the number of training steps/evaluation steps per second performed by the INR model with 1080p inputs/outputs. The model compression MACs and encoding/decoding time are measured by the steps for performing quantization and entropy coding. The complexity figures are calculated based on NVIDIA RTX 4090 GPU with FP16.
@@
The complexity figures of NVRC with the UVG dataset mercat2020uvg are provided in Table 3. When compared to the original HiNeRV kwan2023hinerv, the proposed method (with HiNeRV as its INR network) is associated with increased computational complexity. However, the MACs figure is still significantly lower than that of other learning-based video codecs (e.g., DCVC-DC li2023neural), which allows faster decoding. It should be noted that the complexity figures shown here are obtained based on research source code that has not been optimized for latency. The actual latency of INR and entropy coding can be further reduced by (1) optimizing the implementation of the INR and entropy models, (2) performing lower precision computation, and (3) implementing parallel decoding between different resolution feature grids.
@@ 4.4 Ablation study
Table 4: Ablation studies on the UVG dataset mercat2020uvg. Results are BD-rates.
Metric NVRC (V1) (V2) (V3) (V4) (V5)
PSNR 0.00% 13.04% 11.06% 23.37% 30.84% 14.42%
MS-SSIM 0.00% 13.84% 10.24% 23.88% 30.07% 8.54%
To evaluate the contribution of the main components in NVRC, an ablation study was performed based on the UVG dataset mercat2020uvg, using the configurations in Section 4.1, but four rate points for each variant.
Alternative entropy model settings. We compared different combinations of entropy models for encoding feature grids $\theta_{grid}$ and network parameters $\theta_{layer}$: Context model + dual-axis conditional Gaussian model (Default setting in NRVC), (V1) Context model + per-tensor conditional Gaussian model, (V2) per-tensor conditional Gaussian model + dual-axis conditional Gaussian model, (V3) per-tensor conditional Gaussian model + per-tensor conditional Gaussian model.
@@
Learned quantization steps. We also compared the use of learned quantization steps $\phi_{quant}$ and $\psi_{quant}$ (Default setting in NRVC) and (V5), a new variant with fixed quantization steps for grids, where the log-step size is set to 4-4.
@@
In addition, we conducted experiments to evaluate the effects of fully end-to-end optimization and alternating optimization on selected challenging sequences from the UVG dataset (Jockey and ReadySetGo) mercat2020uvg. When removing fully end-to-end optimization, the variant without rate loss in the first stage exhibits up to a 35% BD-rate increase compared to the proposed model. However, this loss diminishes if the number of epochs in the second stage increases. With the proposed alternating optimization, we did not observe any noticeable difference in performance. Nevertheless, without alternating optimization, the training step time can increase by up to 40% under our experimental settings.